OpenStudy (logan13):

A cylindrical cardboard mailing tube has a diameter of 90 mm. and a height of 600 mm. A total of 20% of the tube’s capacity is filled with spherical packing foam. Each foam sphere has a radius of 15 mm. Approximately how many pieces of foam are in the tube? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (deadpoolroxz):

give me a sec dude

OpenStudy (deadpoolroxz):

Calulate Cylinder Volume: remember; Vcyl = 3.14(r)2h Vcyl = 3.14 (45mm)2 600 mm Vcyl = 3.14 (2025mm2) 600mm Vcyl = 3815100 mm3 or 3.8 x106 mm3 Next calculate the percentage of volume taken up by styrofoam balls 3815100 x 20/100 = 763020 Vstyroballvol = 763020mm3 {as a check calculate} 80% 3815100 x 80/100 = 3052080 mm3 20% 3815100 x 20/100 = 763020 mm3 100% = 80% + 20% = 3815100 mm3 Now calculate the Volume of the Sphere: Vsphere = 4/3*3.14*(r)3 Vsp = 4/3*3.14*(15mm)3 Vsp = 4/3*3.14*(3375) Vsp = 14130 mm3 for 1 stryofoam ball Lastly divide the amount of 20% vol of cylinder by one ball volume 763020mm3 / 14130mm3 = 54 styrofoam balls will fill or are in the tube.

OpenStudy (deadpoolroxz):

54 styrofoam balls will fill or are in the tube.

OpenStudy (logan13):

thanks

OpenStudy (jandr):

you could be right, privided we crush the balls first. I see no provision in your computation for the space between the balls...

OpenStudy (logan13):

he was right

OpenStudy (deadpoolroxz):

*smarts*